# Enhancing LLM-Based Recommendations Through Personalized Reasoning

### v1

@@ Abstract.
Current recommendation systems powered by large language models (LLMs) often underutilize their reasoning capabilities due to a lack of explicit logical structuring. To address this limitation, we introduce CoT-Rec, a framework that integrates Chain-of-Thought (CoT) reasoning into LLM-driven recommendations by incorporating two crucial processes: user preference analysis and item perception evaluation. CoT-Rec operates in two key phases: (1) personalized data extraction, where user preferences and item perceptions are identified, and (2) personalized data application, where this information is leveraged to refine recommendations. Our experimental analysis demonstrates that CoT-Rec improves recommendation accuracy by making better use of LLMs’ reasoning potential. The implementation is publicly available at https://anonymous.4open.science/r/CoT-Rec.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 2.1.1. LLM-as-Retriever
LLM-as-Retriever leverages LLMs to recall a set of potentially relevant items from an entire item set based on a user’s interaction history. To ensure that the retrieved items remain within item set, three key paradigms have been proposed: bi-step grounding, indexing, and modal alignment. Bi-Step Grounding (Lin et al., 2024c; Gao et al., 2024) retrieves items by measuring the similarity between the textual output of the LLM and the candidate set. A pioneering work in this area is BIGRec (Bao et al., 2023a), which reformulates the recommendation task by first grounding the LLM’s output from the language space to the recommendation space, and subsequently aligning it with the actual item space. Indexing (Chen et al., 2024b; Li et al., 2024b) discretizes items into semantically meaningful tokens and employs beam search for retrieval. A notable example is LC-Rec (Zheng et al., 2024b), which aligns language tokens with item index tokens through task-specific optimizations, effectively bridging the gap between language representations and item indexing. Modal Alignment (Yu et al., 2024; Li et al., 2023; Chen et al., 2024a; Zhang et al., 2024g) transforms the semantic vectors encoded by the Collaborative Retrieval Model (CRM) to align them with the semantic space of the LLM, replacing the traditional next-token prediction head with a next-item prediction head. This approach seamlessly integrates collaborative filtering information into the LLM, leading to significant improvements in retrieval performance.
@@ 2.1.2. LLM-as-Ranker
The LLM-as-Ranker paradigm (Cao et al., 2024; Luo et al., 2024b) requires LLMs to either rank a set of candidates based on a user’s interaction history (list-wise ranking) or predict the likelihood of user interaction with a specific item (point-wise ranking). Point-wise Ranking. TALLRec (Bao et al., 2023b) represents a pioneering effort in this domain. Subsequent studies have introduced notable advancements, including integration with Click-Through Rate (CTR) models (Lin et al., 2024a, b), optimization of user preference modeling (Zheng et al., 2024a), and improvements in text-like encoding techniques (Zhang et al., 2024a). List-wise Ranking. LLMRank (Hou et al., 2024b) serves as a foundational work in this area. Building on this, later research has explored various extensions, such as interpretable cross-domain recommendation (Petruzzelli et al., 2024), intent-driven session-based recommendation (Sun et al., 2024b), and comprehensive LLM-powered recommendation systems (Kim et al., 2024). Additionally, researchers have sought to enhance ranking performance through fine-tuning strategies (Yue et al., 2023; Liao et al., 2024; Chen et al., 2024c; Liu et al., 2022) and prompt optimization techniques (Wang and Lim, 2024).
@@ 2.2. LLM-enhanced RSs
LLM-enhanced recommender systems (RSs) leverage LLMs to enhance the capabilities of Collaborative Retrieval Models (CRM) during the training phase, while LLMs are not required during inference. Depending on the type of knowledge provided by the LLM, some studies utilize LLMs to construct or optimize graphs that encode structural knowledge for CRM (Hu et al., 2024; Zhang et al., 2024f; Sakurai et al., 2024; Wang et al., 2024a; Yang et al., 2024; Liu et al., 2024a). Others introduce interaction information into CRM by generating synthetic interactions (Wang et al., 2024b; Wei et al., 2024). Additionally, certain works enhance CRM inputs by optimizing features (Jia et al., 2024; Wang et al., 2024c; Liu et al., 2023c) or generating textual content (Zhang et al., 2024d; Du et al., 2024; Sun et al., 2024c; Xi et al., 2024). Furthermore, some approaches improve CRM’s ability to learn high-quality representations by leveraging embeddings (Geng et al., 2024; Wang et al., 2024d; Cui et al., 2024; Liu et al., 2024b; Harte et al., 2023; Zhang et al., 2024e; Ren et al., 2024).
@@ 3. Preliminaries
In this section, we introduce the sequential recommendation task and two methods that we use as backbones.
Sequential Recommendation. Given a user $u$ who has interacted chronologically with an item sequence $[i_{1},i_{2},\dots,i_{n}]$, the goal of sequential recommendation is to predict the next item $i_{n+1}$ that the user is likely to engage with.
@@ 5.1.3. Backbones
We select SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) as the CRM for the retrieval stage, and Qwen2.5-7B-Instruct as the LLM for the ranking stage. The hyperparameters can be found in the code repository. Due to space limitations, we only present the results of SASRec. The BERT4Rec results are available in the repository and lead to consistent conclusions.
@@ 5.2.2. LLM-as-Ranker
Table 2 presents the effectiveness of different candidate set ranking methods. Here, Retriever refers to the retrieval model used. “CRM” refers to the basic SASRec model (as in Table 1 with “User Embedding=None” and “Item Embedding=Random”), while “CRM++” denotes its improved version (as in Table 1 with “User Embedding=Preference” and “Item Embedding=Description”). The Ranker denotes the ranking method. “None” indicates that no LLM-based ranking is applied, and the CRM retrieval results are used directly. “LLM” performs ranking without user preference and item perception. “LLM+” incorporates user preferences and objective item descriptions, while “LLM++” further integrates user preferences and subjective item perception of the candidate set.
The experimental results show that ranking with “LLM”, “LLM+”, and “LLM++” generally leads to a significant improvement in recommendation performance compared to no ranking (“Ranker=None”). The only exception occurs in the Yelp dataset, where “Retriever=CRM” and “Ranker=LLM” perform similarly to the no-ranking setting. Additionally, while “LLM++” and “LLM+” offer only a slight improvement in accuracy over “LLM”, they substantially reduce position bias. Furthermore, “LLM++” achieves greater positional bias reduction than “LLM+” by incorporating subjective perception rather than objective descriptions. This may be because integrating user preferences and personalized perception helps the LLM focus on task-relevant information, thereby mitigating its inherent position bias in ranking.
@@ 6. Conclusions
We propose CoT-Rec, which enhances LLM-powered recommendations via two key CoT processes: user preference analysis and item perception analysis. Experimental results effectively validate our approach. Inference efficiency analysis shows that CoT-Rec introduces negligible overhead, suggesting its potential for industrial applications. Expanding CoT-Rec to other backbones is a promising avenue for future research.

### v2

# Improving LLM-powered Recommendations with Personalized Information

@@ Abstract.
Due to the lack of explicit reasoning modeling, existing LLM-powered recommendations fail to leverage LLMs’ reasoning capabilities effectively. In this paper, we propose a pipeline called CoT-Rec, which integrates two key Chain-of-Thought (CoT) processes—user preference analysis and item perception analysis—into LLM-powered recommendations, thereby enhancing the utilization of LLMs’ reasoning abilities. CoT-Rec consists of two stages: (1) personalized information extraction, where user preferences and item perception are extracted, and (2) personalized information utilization, where this information is incorporated into the LLM-powered recommendation process. Experimental results demonstrate that CoT-Rec shows potential for improving LLM-powered recommendations. The implementation is publicly available at https://github.com/jhliu0807/CoT-Rec.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730211isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Recommender systems
1. Introduction
@@ 2.1.1. LLM-as-Retriever
LLM-as-Retriever leverages LLMs to recall a set of potentially relevant items from an entire item set based on a user’s interaction history. Bi-Step Grounding (Lin et al., 2024c; Gao et al., 2024; Bao et al., 2023a) retrieves items by measuring the similarity between the textual output of the LLM and the candidate set. Indexing (Chen et al., 2024b; Li et al., 2024b; Zheng et al., 2024b) discretizes items into semantically meaningful tokens and employs beam search for retrieval. Modal Alignment (Yu et al., 2024; Li et al., 2023; Chen et al., 2024a; Zhang et al., 2024g) transforms the semantic vectors encoded by the CRM to align them with the semantic space of the LLM, replacing the traditional next-token prediction head with a next-item prediction head.
@@ 2.1.2. LLM-as-Ranker
The LLM-as-Ranker paradigm (Cao et al., 2024; Luo et al., 2024b) requires LLMs to either rank a set of candidates based on a user’s interaction history (list-wise ranking) or predict the likelihood of user interaction with a specific item (point-wise ranking). Point-wise ranking predicts whether a specific item will be liked by a user (Bao et al., 2023b; Lin et al., 2024a, b; Zheng et al., 2024a; Zhang et al., 2024a), whereas list-wise ranking generates a personalized ranking for an entire set of candidate items (Hou et al., 2024b; Petruzzelli et al., 2024; Sun et al., 2024b; Kim et al., 2024; Yue et al., 2023; Liao et al., 2024; Chen et al., 2024c; Liu et al., 2022; Wang and Lim, 2024).
@@ 2.2. LLM-enhanced RSs
LLM-enhanced recommender systems (RSs) leverage LLMs to enhance the capabilities of CRM during the training phase, while LLMs are not required during inference. Depending on the type of knowledge provided by the LLM, some studies utilize LLMs to construct or optimize graphs that encode structural knowledge for CRM (Hu et al., 2024; Zhang et al., 2024f; Sakurai et al., 2024; Wang et al., 2024a; Yang et al., 2024; Liu et al., 2024a). Others introduce interaction information into CRM by generating synthetic interactions (Wang et al., 2024b; Wei et al., 2024). Additionally, certain works enhance CRM inputs by optimizing features (Jia et al., 2024; Wang et al., 2024c; Liu et al., 2023c) or generating textual content (Zhang et al., 2024d; Du et al., 2024; Sun et al., 2024c; Xi et al., 2024). Furthermore, some approaches improve CRM’s ability to learn high-quality representations by leveraging embeddings (Geng et al., 2024; Wang et al., 2024d; Cui et al., 2024; Liu et al., 2024b; Harte et al., 2023; Zhang et al., 2024e; Ren et al., 2024).
@@ 3. Preliminaries
SASRec and LlamaRec are the backbones used in this paper.
@@ 5.1.3. Backbones
We select SASRec as the CRM for the retrieval stage, and Qwen2.5-7B-Instruct as the LLM for the ranking stage.
@@ 5.2.2. LLM-as-Ranker
Table 2 presents the effectiveness of different candidate set ranking methods. Here, Retriever refers to the retrieval model used. “CRM” refers to the basic SASRec model (as in Table 1 with “User Embedding=None” and “Item Embedding=Random”), while “CRM++” denotes its improved version (as in Table 1 with “User Embedding=Preference” and “Item Embedding=Description”). The Ranker denotes the ranking method. “None” indicates that no LLM-based ranking is applied, and the CRM retrieval results are used directly. “LLM” performs ranking without user preference and item perception. “LLM++” incorporates user preferences and subjective item perception.
The experimental results show that ranking with “LLM” and “LLM++” generally leads to a significant improvement in recommendation performance compared to no ranking (“Ranker=None”). Additionally, while “LLM++” offers only a slight improvement in accuracy over “LLM”, it substantially reduce position bias. This may be because integrating user preferences and personalized perception helps the LLM focus on task-relevant information, thereby mitigating its inherent position bias in ranking.
@@ 6. Conclusions
We propose CoT-Rec, which enhances LLM-powered recommendations via two key CoT processes: user preference analysis and item perception analysis. Experimental results effectively validate our approach. Inference efficiency analysis shows that CoT-Rec introduces negligible overhead, suggesting its potential for industrial applications.
@@ Acknowledgements.
This work is supported by National Natural Science Foundation of China (NSFC) under the Grant No. 62372113. Peng Zhang is a faculty of School of Computer Science, Fudan University. Tun Lu is a faculty of School of Computer Science, Shanghai Key Laboratory of Data Science, Fudan Institute on Aging, MOE Laboratory for National Development and Intelligent Governance, and Shanghai Institute of Intelligent Electronics & Systems, Fudan University.